OpenStudy (anonymous):

What are the critical values of f(x) = x-1 / x + 3

OpenStudy (solomonzelman):

find the \(\large\color{black}{ f'(x) }\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it - x-1 / x + 3

OpenStudy (solomonzelman):

that is not the derivative of your function. or are you trying to say that it id the function?

OpenStudy (anonymous):

no I tried to do quotient rule

OpenStudy (anonymous):

isnt the quotient rule f'g - fg' / g^2 ??

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE dy}{\LARGE dx} \left(\begin{matrix} \frac{\LARGE x-1}{\LARGE x+3} \\ \end{matrix}\right)=\frac{\LARGE (x+3)(x-1)'-(x+3)'(x-1)}{\LARGE (x+3)^2} }\) makes sense?

OpenStudy (anonymous):

so then its (x+3)(1) - (1)(x-1) / (x+3)^2 ?

OpenStudy (solomonzelman):

yes, so the top would be?

OpenStudy (anonymous):

wouldny the top be -x-1 because the x+3 divides with one from the denominator

OpenStudy (anonymous):

or does it stay 4 / (x+3)^2

OpenStudy (solomonzelman):

yup 4/(x+3)^2

OpenStudy (anonymous):

why cant you divide the x+3 if there are two of them

OpenStudy (solomonzelman):

\(\large\color{black}{ f'(x)= \frac{\LARGE 4}{\LARGE (x+3)^2} }\)

OpenStudy (solomonzelman):

when you set the \(\large\color{black}{ f'(x)=0 }\) , is there an x value that satisifes the statement?

OpenStudy (solomonzelman):

please, this isn't your question

OpenStudy (solomonzelman):

@ilovereyvis_x3 \(\large\color{black}{ f'(x)= \frac{\LARGE 4}{\LARGE (x+3)^2} }\) \(\large\color{black}{ 0= \frac{\LARGE 4}{\LARGE (x+3)^2} }\) .....is there an x-solution?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

yes, there is no solution. So there are NO CRITICAL NUMBERS.

OpenStudy (solomonzelman):

\(\large\color{black}{ f'(x)= \frac{\LARGE 4}{\LARGE (x+3)^2} }\) would yield that x=-3 is a critical numb. but here since -3 is not part of the domain of \(\large\color{black}{ f(x) }\), it is not a critical numb. (x=-3 would have been a critical numb, if it was in the domain of \(\large\color{black}{ f(x) }\) )

OpenStudy (anonymous):

ok , but can you help me on how the derivative ended up being 4/ (x+3)^2

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE dy}{\LARGE dx} \left(\begin{matrix} \frac{\LARGE x-1}{\LARGE x+3} \\ \end{matrix}\right)=\frac{\LARGE (x+3)(x-1)'-(x+3)'(x-1)}{\LARGE (x+3)^2} }\) this is the quotent rule.

OpenStudy (solomonzelman):

Correct?

OpenStudy (anonymous):

yeah but how is (x+3) - (x-1) = 4???

OpenStudy (solomonzelman):

\(\large\color{black}{ - (x-1) ~~~~~\Rightarrow~~~~~-x+1 }\)

OpenStudy (solomonzelman):

I keep disconnecting, sorry-:(

OpenStudy (anonymous):

it's fine don't worry i do also

OpenStudy (solomonzelman):

Again, sjut now

OpenStudy (solomonzelman):

Anyway, see how: \(\large\color{red}{ - (x-1) ~~~~=~~~~-x+1 }\) ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

so, \(\large\color{red}{ (x+3) - (x-1) = 4 }\) \(\large\color{red}{ ^{\Huge \color{white}{|}} x+3 - x+1 = 4 }\)

OpenStudy (solomonzelman):

the \(\large\color{red}{X}\)s cancel, and you get 4.

OpenStudy (anonymous):

OH !! ok I understand thank you !!!!

OpenStudy (solomonzelman):

anytime:)